OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that contains the following points? (-2,-20),(0,-4),(4,-20) I tried plugging them into the standard equation and using elimination, but I got no solution. I have no idea how to go about this problem.

OpenStudy (hlambach):

Hum, let me try doing that and see if I get the answer.

OpenStudy (hlambach):

In standard form (y=ax^2+bx+c), c is the y-intercept. One of the points gives the y intercept. Start there.

OpenStudy (anonymous):

c=-4

OpenStudy (hlambach):

Yep.

OpenStudy (anonymous):

I get to -20=-2a^2 -2b -4 and -20=4a^2+4b-4 and I don't know what to do next

OpenStudy (anonymous):

4a for the first one and 16a for the second once simplified

OpenStudy (hlambach):

So use the second and third equation and use the elimination thing.

OpenStudy (hlambach):

Sorry, that meant to send way earlier.

OpenStudy (anonymous):

I keep getting -32=8a and a=0

OpenStudy (anonymous):

Which does not make sense...

OpenStudy (hlambach):

Anyways...OH, I know what you did wrong! You are squaring the a, not the x. :)

OpenStudy (anonymous):

... so what do I need to do?

OpenStudy (hlambach):

standard form is y=ax^2+bx+c So the two formulas would be \[y=a(-2)^{2}+b(-2)+c\] \[-20=a(4)^{2}+b(4)-4\]

OpenStudy (hlambach):

Now try.

OpenStudy (anonymous):

Is the first equation supposed to be "y=" or -20?

OpenStudy (hlambach):

Sorry, the y is supposed to be -20.

OpenStudy (hlambach):

My bad. :)

OpenStudy (anonymous):

4a+(-2b)-4=-20 -16=4a-2b -20= 16a+4b-4 -16=16a+4b

OpenStudy (hlambach):

Yup! Now do elimination

OpenStudy (anonymous):

-32=8a-4b -16=16a+4b -48=24a

OpenStudy (anonymous):

-2=a

OpenStudy (hlambach):

There you go! Glad I could help! I've done that mistake before. I was as lost as you. :)

OpenStudy (anonymous):

-16=-2(16)+4b -16=-32+4b 16=4b b=4

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (hlambach):

No problem! You can best show your thanks by giving out a medal. :)

OpenStudy (hlambach):

Thank you! If you need any help, you can just tag me. :)

OpenStudy (anonymous):

When I make you a best response, does that give a medal?

OpenStudy (hlambach):

Yes! Thanks again! :)

OpenStudy (anonymous):

No, thank you (;